OpenStudy (anonymous):

If x varies inversely with y and x = 8 when y = 6, find y when x = 10.

undeadknight26 (undeadknight26):

hmm well lets say we multiply x by 2 so 10 x 2 = 20 and divide 6 by 2 making it 6/2=3. i believe that is how its done..not sure.

OpenStudy (anonymous):

A. y = 5/12 B. y = 4.8 C. y = 7.5 D. y = 40/3

OpenStudy (anonymous):

those are the answers

OpenStudy (anonymous):

or one of them is

undeadknight26 (undeadknight26):

*tilts head to the side* uh...i dunno now XD @agent0smith @eliassaab

OpenStudy (anonymous):

me either :(

OpenStudy (anonymous):

@AngelWilliams16 @thadyoung @tester97

OpenStudy (tester97):

Ok first off do you know what it means when you have indirect variation?

OpenStudy (thadyoung):

@bibby @mathmale

OpenStudy (tester97):

How do you expect help if you dont tell me what you know? >.<

OpenStudy (anonymous):

@Cookie_009

OpenStudy (agent0smith):

x varies inversely with y means x = k/y where k is some constant. First, find it by plugging in x = 8 when y = 6

OpenStudy (anonymous):

Well my sister was helping me and she told me it was A but i thought it was C

OpenStudy (agent0smith):

y = k/x where k is some constant. First, find it by plugging in x = 8 and y = 6

OpenStudy (agent0smith):

Then once you have k, plug it in and x=10 into: y = k/x and find y.

OpenStudy (anonymous):

so i was right its C

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

I don't know, haven't worked it out. Show your work of how you got the answer, and I or someone will check it.

OpenStudy (anonymous):

ughhh i just need to know it im just checking my asnwers

OpenStudy (tester97):

k y= ------ x so plug in 8 and 6 k 6= ------ 8 multiply 8 on both sides and you have 48=k then plug k back in when x=10 48 y= ------ 10 do simple division y= 4.8 so the answer would be choice B

OpenStudy (anonymous):

thankyou can u help me with 1 more!!

OpenStudy (agent0smith):

^if you'd showed your work, we might've noticed what you were doing wrong, which would help you.